OpenStudy (anonymous):

What is the linear speed of a point a) on the equator. b) On the Arctic Circle (latitude 66.5 degrees N), and c) at a latitude of 45.0 degrees N, due to the Earth's rotation. Im not sure how to start this!! :( For part A in my book it says the Earths radius is 1.74x10^3 km. Would I convert this to meters and use it in this equation?: (1740m)(360/2pi rad)? Am I on the right track?

OpenStudy (experimentx):

 Use the relation \[\huge V = r cross \omega=r\omega \sin\theta\] = Read \( \theta \) from top angle.

OpenStudy (experimentx):